Please solve 7w/4 + 5 = 3w/10 + 1

Answer by sham.alex@hotmail.com(1)

7w/4 + 5 = 3w/10 + 1
or, 7w/4 - 3w/10 = 1-5
or, (35w-6w)/20 = -4 ,( in L.H.S. taking lcm of 4 and 10).
or, 29w = -80
or, w = -80/29
= -2.76 ,(taking two decimal places)